OpenStudy (anonymous):

find the antiderivative of : 1.4e^(0.07t) dt

OpenStudy (anonymous):

rawr. integrate it!. \[\int\limits of (that thing) = 1.4 \int\limits (e^0.07t)dt. u = 0.07t, du/dt = 0.07. \therefore dt = du/0.07. and. then do 1.4/0.07 integrate e^u. and get 1.4 e^u, sub for u and get 1.4/0.07 e^0.07t\]

OpenStudy (anonymous):

sigh this hates me. .. sub for u. then get (1.4/0.07) integral of e^u, then by integrating you get the integral of e^u is e^u. then get (1.4/0.07)e^(0.07t) as your answer